OpenStudy (anonymous):

how to solve laplace of sin^3 2t?

OpenStudy (sidsiddhartha):

use sin(x)^3 = (3/4)sin(x) - (1/4)sin(3x) or u can use euler formula \[\sin(at)=\frac{ e ^{iat} -e ^{-iat}}{ 2i}\]

OpenStudy (anonymous):

thnx

OpenStudy (sidsiddhartha):

ur wlcm :)